The perimeter of a rectangle is 146 inches. The length exceeds the width by 29 inches. Find the length and the width.
Answer by msgmhenderson(11)

The perimeter of a rectangle is 146 inches. The length exceeds the width by 29 inches. Find the length and the width.
perimeter formula for rectangle is 2W * 2L where W represent width and L represent length
We know that the length is L + 29 for one side so 2 sides is 2(L + 29)
and the width is W so for 2 sides is 2W
2 (L + 29) + 2W = 146
2L +(2)*(29) + 2W = 146
2L + 58 +2W = 146
2L + 2W = 146 - 58
2L + 2W = 88
88/4 WHICH IS THE FOUR SIDES
22 IS WIDTH
22 + 29 IS LENGTH
51 IS LENGTH